OpenStudy (anonymous):

A boat can travel 45miles on 7 gallons of gasoline. How much gasoline will it need to go 78 miles? * You would the proportion as 45/7=78/x true False

OpenStudy (texaschic101):

TRUE miles/gallons : 45/7 = 78/x....45 miles to 7 gallons = 78 miles to x gallons